# Quantum reservoir computing maps data onto the Krylov space

### v1

@@ Abstract
This results in the spaces $\mathcal{F}_{k}$ being linearly independent and ensures that each observable $O_{k}$ is mapped onto a space $\mathcal{F}_{k}$. In the next step, we use these spaces to define Krylov observability, a measure to capture the phase space dimension of various operators.
Proposition 1 (Krylov Observability).
$\mathcal{G}_{k}=\{O_{k}(\tau_{1}),O_{k}(\tau_{2}),\dots,O_{k}(\tau_{M_{k}})\}$
$p_{k}(T)=1+\sum_{j=1}^{R_{k}-1}(1-f(O_{k}(\tau_{j}),O_{k}(\tau_{j+1}))),$ (5)
where $f$ is a function with $f(A,B)=1$ if $A=B$. In this study, we set
$|\displaystyle f(A,B):=\mathrm{F}(A,B)=\absolutevalue{\mathrm{Tr}(\frac{A^{\dagger}}{\norm{A}}\frac{B}{\norm{B}}}$ | (6)
$\mathcal{O}_{K}(T)=\sum_{k=1}^{K}p_{k}(T).$ (7)
$H_{I}=\sum_{i=1,j>i}^{N_{S}}J_{ij}X_{i}X_{j}+\sum_{i=1}^{N_{S}}hZ_{i}.$ (8)
$\partial_{t}\rho=-i[H,\rho].$ (9)
[…]
Reservoir computing research often makes use of the information processing capacity ($\mathrm{IPC}$) as an expressivity measure. The $\mathrm{IPC}$ evaluates how well prior inputs and combinations of prior inputs can be reconstructed non-linearly [65]. The first-order $\mathrm{IPC}_{1}$, also called memory capacity, measures how well the system can reconstruct prior data [67]. The definition for computing the $\mathrm{IPC}$ is given in [65, 61, 68, 69, 70]. The sum over all $\mathrm{IPC}$s is called the total $\mathrm{IPC}$, or simply $\mathrm{IPC}$, i.e., $\mathrm{IPC}=\sum_{i}\mathrm{IPC}_{i}$. [71, 72] provides analytical evidence of the influence of eigenvalues on the $\mathrm{IPC}$. Furthermore, it has been shown that $\mathrm{IPC}$ can help predict task performance in reservoir computing [73, 74], and there is a typically observed non-linearity-memory trade-off, which was observed in quantum reservoirs as well [61].
Results—To analyze the ability of our Krylov observability measure $\mathcal{O}_{K}$ to quantify the computational capabilities of a quantum system, we compare it with the $\mathrm{IPC}$. We consider ten different Ising Hamiltonians, each with four sites and uniformly distributed coupling constants $J_{i,j}$ on the interval $[0.25,0.75]$. For each Hamiltonian, we consider a quantum reservoir where the first site ($Z_{1}$), the first two sites ($Z_{1,2}$), and all four sites ($Z_{1,2,3,4}$) are measured in the Pauli-$z$ direction. We compute the information processing capacity $\mathrm{IPC}$ and the Krylov observability $\mathcal{O}_{K}$ for various clock cycle lengths $T$ and for various time-multiplexing measurements $V$.
Figure 1: Observability $\mathcal{O}_{K}$ and $\mathrm{IPC}$ in dependence on multiplexing $V$ and clock cycle $T$.
Figure 1 shows the results for the averaged $\mathrm{IPC}$ (left column) and the averaged $\mathcal{O}_{K}$ (right column). In each plot, the $y$-axis represents the number of measurements for the corresponding measurement scheme and the $x$-axis represents the clock cycle $T$. The first row shows results when only the first site is measured $Z_{1}$.
As the clock cycle $T$ and the number of measurements $V$ increase, so does the phase space dimension, i.e., the Krylov observability $\mathcal{O}_{K}$, onto which the data is mapped. At $T_{\mathrm{sat}}(Z_{1})\approx 36$ and $V_{\mathrm{sat}}(Z_{1})\approx 110$, the full Krylov space is explored, as indicated by the saturation of $\mathcal{O}_{K}$. Further increases in the number of measurements and clock cycles result only in measurements that are linearly dependent on all previous ones, thus not further increasing the expressivity of the reservoir. This is observed in the almost identical behavior of $\mathrm{IPC}$ and $\mathcal{O}_{K}$. To better quantify the behavior of the Krylov observability $\mathcal{O}_{K}$ and information processing capacity $\mathrm{IPC}$, we compute the Pearson correlation factor between them. In the case where only one observable is measured, a Pearson correlation factor of $P_{C}=0.97$ is reached. Measuring the first two sites yields results in the second row, where both measures show similar behavior with slightly smaller values for $\mathcal{O}_{K}$. The number of virtual nodes at which $\mathrm{IPC}$ and $\mathcal{O}_{K}$ saturate reduces to $T_{\mathrm{sat}}(Z_{1,2})\approx 30$ and $V_{\mathrm{sat}}(Z_{1,2})\approx 95$ with $P_{C}=0.97$. Measuring all four sites moves the saturation to smaller values, $V_{\mathrm{sat}}(Z_{1,2,3,4})\approx 75$ and $T_{\mathrm{sat}}(Z_{1,2})\approx 18$. The earlier saturation when measuring multiple observables is due to the higher number of measurements $N_{R}$ as the total number of measurements is given by $N_{R}=VK$. Additionally, the space dimension of $\mathcal{F}(Z_{1})$ is smaller than that of $\mathcal{F}(Z_{1},Z_{2})$ and $\mathcal{F}(Z_{1},Z_{2},Z_{3},Z_{4})$. One interesting observation, particularly for $\mathcal{O}_{K}$ in the third row, is oscillations over time, arising from the observability measure construction, where the fidelity of time-evolved observables is computed. To better understand this behavior, Fig. 2 shows the $\mathrm{IPC}$ (blue) and $\mathcal{O}_{K}$ (red) as functions of the clock cycle $T$ for a fixed number of measurements $V=30$ (left), and as functions of $V$ for a fixed clock cycle $T=20$ (right). In the left plot, $\mathcal{O}_{K}$ increases slightly later but reaches a maximum with oscillations, while the $\mathrm{IPC}$ reaches a maximum around $T\approx 20$ and remains stable. The standard deviation is shown as a fill in the corresponding color and is small for both systems. The oscillation is due to the small number of qubits, as quasiperiodic behavior is more apparent [40]. The right plot shows the dependence on the number of measurements $V$ for constant $T=20$. The behavior of $\mathcal{O}_{K}$ and $\mathrm{IPC}$ is, except for an offset, almost identical.
Figure 2: Observability $\mathcal{O}_{K}$ and $\mathrm{IPC}$ in dependence on multiplexing $V$ and clock cycle $T$.
Conclusion and Discussion— This work showed that instead of computing different powers of the Liouvillian, time-evolved operators can be used to construct an equivalent space in Theorem 1. Building upon this, we have shown how observability spaces can be constructed that consider the possible dependence of both spaces, and an algorithm for such a computation is provided. This enabled our work to propose Krylov observability as a measure to capture the phase space dimension of evolved observables, see Proposition 1. Krylov observability was then tested in the framework of quantum reservoir computing. We showed that the information processing capacity, consisting of thousands of quantum reservoir tasks, and the Krylov observability exhibit almost identical behavior, with correlation factors of $P_{C}=0.97$. The high correlation between the information processing capacity and Krylov observability suggests that quantum reservoir computing can be interpreted as a map of classical data onto the Krylov spaces discussed in this work. Second, the field of Krylov complexity is validated as an effective method for understanding quantum systems through Krylov observability, which can predict how well macroscopic information (i.e., encoded input series) can be reconstructed and non-linearly transformed in a quantum system. This study further demonstrates that information processing capacity, or macroscopic information, might be an interesting technique to describe the physics of quantum systems. Therefore, quantum reservoir computing might become a tool to further the understanding of quantum systems. Lastly, we show that Krylov observability serves as an expressivity measure that is fast to compute, with a simulation time for one Hamiltonian of 30 seconds, while the corresponding computation of the information processing capacity can take over 150 hours. Programming information processing capacity also requires careful consideration of combinatorics and suitable cutoffs, whereas programming Krylov observability is straightforward. Lastly, Krylov observability can easily adapt as an expressivity measure for variational quantum machine learning and quantum extreme learning.

### v2

@@ Appendix A Algorithm for the construction of disjiont Krylov spaces
Appendix B Quantum Zeno Time
Let us consider the fidelity overlap between the observable $O$ and the time-evolved observable at $O(t)$ given by
$F(O,O(t))=\absolutevalue{\mathrm{Tr}\Big{(}\frac{O^{\dagger}O(t)}{\norm{O}\norm{O(t)}}\Big{)}}$ | (11)
We normalize $O:=O/\norm{O}$. The time-evolved operator up to second order can be approximated as
$O(t)=e^{iHt}Oe^{-iHt}=O+it[H,O]-\frac{t^{2}}{2}[H,[H,O]]+\mathcal{O}(t^{3})$ (12)
This leads to the overlap of
$\mathrm{F}(t):=\mathrm{Tr}(O^{\dagger}O(t))=1+itB-\frac{t^{2}}{2}C+\mathcal{O}(t^{3})$ (13)
where:
$B=\mathrm{Tr}(O^{\dagger}[H,O]),\qquad C=\mathrm{Tr}(O^{\dagger}[H,[H,O]])$ (14)
Next we can compute the overlap squarred
$\left|\mathrm{F}(t)\right|^{2}$ $=\left|1+itB-\frac{t^{2}}{2}C\right|^{2}+\mathcal{O}(t^{3})$
$=\left(1+itB-\frac{t^{2}}{2}C\right)\left(1-itB^{*}-\frac{t^{2}}{2}C^{*}\right)+\mathcal{O}(t^{3})$
$=1+2t\,\mathrm{Im}(B)+t^{2}\left(|B|^{2}-\mathrm{Re}(C)\right)+\mathcal{O}(t^{3})$ (15)
In our case it holds that $H=H^{\dagger}$ and $O=O^{\dagger}$ resulting in $C$ being real. We further make use of $O$ and $H$ being real representations, since they only consists of real pauli-matrices. This results in $B=Tr(O[H,O])$ and $C=\mathrm{Tr}(O[H,[H,O]])$ being real as well, such that the overlap is given by
|$\absolutevalue{\mathrm{F}(t)}^{2}$ $=1-t^{2}(C-B^{2})$
$=1-t^{2}(\mathrm{Tr}(O[H,[H,O]])-\mathrm{Tr}(O[H,O])^{2})+\mathcal{O}(t^{3}).$ (16)
With this, we can define the zeno-like time
$\tau_{z}^{-2}=\mathrm{Tr}(O[H,[H,O]])-\mathrm{Tr}(O[H,O])^{2}$ (17)
, which results in
$\left|\mathrm{F}(t)\right|^{2}=1-\frac{t^{2}}{\tau_{z}^{2}}+\mathcal{O}(t^{3}).$ (18)
If we use the Liouvillian, this timescale is given by
$\tau_{z}^{-2}=\mathrm{Tr}(O\mathcal{L}^{2})-\left[\mathrm{Tr}(O\mathcal{L})\right]^{2}$ (19)
and can be interpreted as an expectation value with respect to the operator $O$, i.e., $\langle\mathcal{L}\rangle=\mathrm{Tr}(O\mathcal{L})$. This leads to
$\tau_{z}^{-2}=\langle\mathcal{L}^{2}\rangle-\langle\mathcal{L}\rangle^{2},$ (20)
which closely mirrors the quantum Zeno time for state evolution,
$\tau_{z,(\ket{\psi})}^{-2}=\langle H^{2}\rangle-\langle H\rangle^{2},$ (21)
as discussed in [76, 77], justifying the interpretation as a quantum Zeno time for operator evolution.